# Finite Sample Frequency Domain Identification

### v1

@@ 1 Introduction
Our paper is related to non-parametric system identification, which includes works on both time [20] and frequency domain [21]. Using Gaussian Processes as in [21], where the unknown frequency response follows a Gaussian prior, we can also obtain finite sample guarantees. Here, we follow a different approach and we do not consider Gaussian priors. Note that in this work we focus on qualitative data-independent bounds linking sample requirements to system theoretic properties. Application-oriented, data-dependent bounds have also been studied before [22].
Notation. Let $(H,\mathbb{F},\langle\cdot,\cdot\rangle_{H})$ be a Hilbert space with field $\mathbb{F}=$ or $\mathbb{C}$ and inner product $\langle\cdot,\cdot\rangle_{H}$. For any vector $x\in H$, let $\|x\|\triangleq\sqrt{\langle x,x\rangle_{H}}$ denote the inner product norm. Let $H,V$ be Hilbert spaces with $\mathbb{F}=$ or $\mathbb{C}$ and let $\mathcal{A}:H\rightarrow V$ be any linear map. Let $\|\mathcal{A}\|_{\mathrm{op}}\triangleq\sup_{\|x\|=1}\|\mathcal{A}(x)\|$ denote the operator norm and $\mathcal{A}^{*}$ denote the adjoint operator. If $\left\{b_{i}\right\}_{i\in\mathcal{I}}$ is an orthonormal basis, the Hilbert-Schmidt or Frobenius norm is defined as $\|\mathcal{A}\|_{\mathrm{F}}^{2}\triangleq\sum_{i\in\mathcal{I}}\|\mathcal{A}(b_{i})\|^{2}$. Let $\mathcal{L}_{2}(p)\triangleq\{\left\{x_{k}\right\}^{\infty}_{k=1}:x_{k}\in^{p},\,\sum_{k=1}^{\infty}\|x_{k}\|_{2}^{2}<\infty\}$ be the Hilbert space of $p$-dimensional square summable sequences. A universal constant is a constant that is independent of the problem at hand, e.g., the system or the algorithm. For any integer $M$, let $[M]\triangleq{0,\dots,M-1}$. The $\mathcal{H}_{\infty}$ norm of $G(e^{j\omega}):[0,2\pi)\rightarrow\mathbb{C}^{d_{1}\times d_{2}}$ is given by $\sup_{\omega\in[0,2\pi)}\|G(\omega)\|_{\mathrm{op}}$; it is denoted by $\|G\|_{\mathcal{H}_{\infty}}$.
@@ 2 Problem formulation
where $h_{t}\in\mathbb{R}^{d_{{y}}\times d_{{e}}}$ are the unknown filter coefficients. Let $e_{t}\in\mathbb{R}^{d_{{e}}}$ be i.i.d. zero mean, with covariance $\mathbb{E}{e_{t}e_{t}^{\top}}=\sigma_{e}^{2}I_{d_{{e}}}$, and $K^{2}$-sub-Gaussian, i.e., for any $\xi\in^{d_{{e}}}$
@@
$\left\|u_{t}\right\|_{2}\leq D_{{u}},\,\text{ for all }t\in\mathbb{Z}$
@@ 2.1 Empirical Transfer Function Estimate
The goal of non-parametric frequency domain identification is to estimate the frequency response $G(e^{j\omega})$, given input-output data. We assume access to $d_{{u}}$ experiments of length $N$, that is, data $(u^{(i)}_{0},y^{(i)}_{0},\dots,u^{(i)}_{N-1},y^{(i)}_{N-1})$, for $i=1,\dots,d_{{u}}$. This brings the total number of samples to $N_{\mathrm{tot}}\triangleq d_{{u}}N$. For simplicity, we assume that the trajectories are statistically independent. Our analysis can be extended to the single trajectory case, where we gather all samples sequentially.
@@ 2.2 Excitation Method
for some $0<\sigma^{2}_{u,\ell}$ such that $\sum_{\ell=0}^{M-1}\sigma^{2}_{u,\ell}\leq MD_{{u}}^{2}$, where $D_{{u}}$ is the input upper bound of Assumption 2.
@@ 2.3 Objective
Fix a frequency resolution $M<N$ such that $M$ divides $N$ and denote their ratio by $N_{p}=N/M$. Consider $d_{{u}}$ independent input-ouput trajectories of length $N$$\{u^{(i)}_{t},y^{(i)}_{t}\}_{t\in[N]}$, for $i=1,\dots,d_{{u}}$, generated by system (1) with excitation inputs as in Assumption 3. Determine $\epsilon_{\ell}>0$, $\ell\in[M]$ and $\delta>0$ such that
@@ 3 Finite-sample guarantees for the ETFE
In this section, we focus on estimating the frequency response at the selected frequencies $\{2\pi\ell/M\}_{\ell\in[M]}$. Following the convention of (4), we define the stacked DFTs of the noises and the noiseless outputs as
@@
The first term of the right-hand side captures the transient error $T_{k,N}U^{-1}_{k}$, while the second one captures the error $V_{k}U^{-1}_{k}$ due to stochastic noise. Recall that the total number of samples is equal to $N_{\mathrm{tot}}=d_{{u}}N$. As we increase the number of samples $N_{\mathrm{tot}}$ while keeping $M$ constant, the former term decays at a faster rate of $1/N_{\mathrm{tot}}$ compared to the latter’s $1/\sqrt{N_{\mathrm{tot}}}$. Hence, the non-asymptotic rate is
@@
The rate is similar to the ones for non-asymptotic parametric identification in time-domain [17]; the optimal rate in that line of work is typically of the order $\sqrt{d}$ for some $d$ scaling with the number of unknown parameters. Here, we have a similar scaling of $\sqrt{M}(\sqrt{d_{{u}}}+\sqrt{d_{{y}}})$ times an additional $\sqrt{d_{{u}}}$ dimensional dependence. This is an artifact of imposing a strict input norm bound in Assumption 2. If we allow $\sigma_{{u}},D_{{u}}$ to scale with $\sqrt{d_{{u}}}$ (as is the case for white-noise inputs in the time-domain [17]), we can remove this extra term.
@@
Lemma 1 (Stochastic Transient [1]).
Denote the power spectrum of the noise at frequency $\omega_{k}=2\pi k/N$, for some $k=0,\dots,N-1$, by $\Phi_{v}(k)\triangleq\sum_{t=-\infty}^{\infty}R_{t}e^{-j\omega_{k}t}$, with $R_{-t}=R_{t}^{\top}$. We have
$\|\Phi_{v}(k)-\Phi_{v,N}(k)\|_{\mathrm{op}}\leq 2\|R\|_{\star}/N.$
@@ 3.1 Deterministic transient
Lemma 2 (Deterministic Transient [1]).
@@
$\sup_{k=0,\dots,N-1}\|T_{k,N}\|_{\mathrm{op}}\leq 2\|G\|_{\star}D_{{u}}/\sqrt{N}.$
@@ 3.3 Noise concentration
Under Assumption 1, there exists a universal constant $c$ such that for any $s>0$
$\mathbb{P}\left\{\left|\|V_{k}\|_{\mathrm{op}}-\sqrt{\mathrm{tr}(\Phi_{v,N}(k))}\right|>2s\sqrt{\mathrm{tr}(\Phi_{v,N}(k))}\right\}\leq 9^{2d_{{u}}}2\exp\left(-cs^{2}\frac{\sigma_{{e}}^{4}}{K^{4}}\frac{\mathrm{tr}(\Phi_{v,N}(k))}{\|\Phi_{v,N}(k)\|_{\mathrm{op}}}\right).$ (14)
@@ 6 Conclusion and Future Work
We provide finite-sample guarantees for the ETFE over a selected frequency grid, in the case of open-loop periodic excitation and under strict stability assumptions. By tuning the frequency resolution and exploiting Lipschitz continuity, we also obtain estimation guarantees in the $\mathcal{H}_{\infty}$ norm. An interesting direction for future work is studying finite-sample non-parametric least squares [5, 16] in the frequency domain. This approach could lead to interesting connections between function class complexity and experiment design. Moreover, adding more structure, beyond Lipschitz continuity, will lead to faster rates. Other topics that are left for future work include the estimation of the noise statistics, the prediction problem, and obtaining finite-sample guarantees when we use random inputs instead of deterministic ones. Finally, extending minimax lower bounds to this setting is also open [26].
@@ Appendix A Proof of Theorem 2
$\begin{bmatrix}\mathcal{A}_{11}&\mathcal{A}_{12}&\cdots\\ \mathcal{A}_{21}&\mathcal{A}_{22}&\cdots\\ \vdots\\ \mathcal{A}_{d1}&\mathcal{A}_{d2}&\cdots\end{bmatrix},$
@@
$\|\mathcal{A}\|_{\mathrm{F}}^{2}\triangleq\sum_{i=1}^{d}\sum_{k=0}^{\infty}\left\|\mathcal{A}_{ij}\right\|^{2}.$
@@
Hence, by the Kolmogorov maximal inequality [27, Th. 2.5.3], the series $\lim_{k\rightarrow\infty}\sum_{j=1}^{k}{\mathcal{A}_{ij}z_{j}}$ converges almost surely for any $i=1,\dots,d$ and, thus, $\mathcal{A}(z)$ is almost surely well-defined. ∎
@@
We will consider finite-dimensional truncations of the quadratic form $\|\mathcal{A}(z)\|^{2}_{2}$, apply the result of [17], and take limits. Let $\hat{\mathcal{A}}^{(k)}:\mathcal{L}_{2}(p)\rightarrow^{d}$ be the truncated map with semi-infinite matrix representation
@@
Step a): truncation. First, we will establish the desired inequality for all $X_{k}$. Notice that $\|\hat{\mathcal{A}}^{(k)}(z)\|^{2}_{2}$ is a finite dimensional quadratic form. Hence, Proposition A.1 of [17] applies and provides the bound
@@
$|\|\mathcal{A}(z)\|^{2}-\|\hat{\mathcal{A}}^{(k)}(z)\|_{2}^{2}|\leq 2\left|\left(\hat{\mathcal{A}}^{(k)}(z)\right)^{\top}(\mathcal{A}-\hat{\mathcal{A}}^{k})(z)\right|+\|(\mathcal{A}-\hat{\mathcal{A}}^{k})(z)\|^{2}.$
Let $C_{z}=\sup_{t\geq 1}\mathbb{E}z^{\top}_{t}z_{t}$; due to sub-Gaussianity $C_{z}=cpK^{2}$ for some universal constant $c$–see proof of Lemma 5. Invoking the Cauchy-Schwartz inequality
$\mathbb{E}|\|\mathcal{A}(z)\|^{2}-\|\hat{\mathcal{A}}^{(k)}(z)\|_{2}^{2}|\leq 2C_{z}\|\mathcal{A}\|_{\mathrm{F}}\|\mathcal{A}-\hat{A}^{(k)}\|_{\mathrm{F}}+C_{z}\|\mathcal{A}-\hat{A}^{(k)}\|_{\mathrm{F}}^{2}\rightarrow 0,$
@@
The proof of Theorem 2 now follows from Lemma 6 and the Chernoff bound method. The remaining details are omitted since they are identical to [17, proof of Th. 2.1 ].
@@ Appendix B Proof of Lemma 4
Due to the independence between the experiments
$\mathbb{E}V^{*}_{k}V_{k}=\mathrm{diag}(\mathbb{E}(V^{(i)})^{*}_{k}V^{(i)}_{k})=\mathrm{tr}(\Phi_{v,N}(k))I_{d_{{u}}}.$
@@
Step a: fixed $\xi$. Observe that $V^{(i)}_{k}=\mathcal{F}_{k}((h*e^{(i)}))$ is the output of a linear map applied to the infinite sequence $e^{(i)}\triangleq\{e^{(i)}_{N-1},e^{(i)}_{N-2},\dots\}$, where * denotes convolution. We can rewrite the above linear relation using the notation $V^{(i)}_{k}=H_{k}(e^{(i)})$, where $H_{k}$ represents the map $\mathcal{F}_{k}\circ((h*\cdot))$. For convenience, let us group all noise sequences $e^{(i)}$ into one single sequence $e$. We use the enumeration
@@
$\Xi=\begin{bmatrix}\xi^{*}&0&0&\cdots\\ 0&\xi^{*}&0&\cdots\\ 0&0&\xi^{*}\\ \vdots&&&\ddots\end{bmatrix}.$
Since $\xi^{*}\xi=1$, $\Xi\Xi^{*}$ is the identity operator and has unit norm. Since $V_{k}\xi\in\mathbb{C}^{d_{{y}}}$ is complex-valued, we need to embed the vector to a real vector space to apply the Hanson-Wright inequality. Let $\Re(H_{k}\circ\Xi)$ denote the real part of the operator $H_{k}\circ\Xi$ with $\Im(H_{k}\circ\Xi)$ the respective imaginary part. Then, lift the space to ${}^{2d_{{y}}}$
$\hat{V}_{k}\xi=\mathcal{A}(e)\triangleq\begin{bmatrix}\Re(H_{k}\circ\Xi)(e)\\ \Im(H_{k}\circ\Xi)(e)\end{bmatrix}\in^{2d_{{y}}}.$
Since the sequence $e$ is real-valued
$\xi^{*}\hat{V}^{*}_{k}\hat{V}_{k}\xi=\xi^{*}V^{*}_{k}V_{k}\xi.$
Hence, we have
$\sigma_{{e}}^{2}\|\mathcal{A}\|_{\mathrm{F}}^{2}=\mathbb{E}\|\mathcal{A}(e)\|_{2}^{2}=\mathbb{E}\xi^{\top}V^{*}_{k}V_{k}\xi=\mathrm{tr}(\Phi_{v,N}(k)).$
For the operator norm, we use the bound $\|\mathcal{A}\|_{\mathrm{op}}\leq\|H_{k}\|_{\mathrm{op}}=\sqrt{\|\Phi_{v,N}(k)\|_{\mathrm{op}}}/\sigma_{{e}}$, where the first inequality follows from the fact that $\Xi$ has unit operator norm. The equality follows from the fact that $\Phi_{v,N}(k)=\mathbb{E}(V^{(i)}_{k}(V^{(i)}_{k})^{*})=\sigma_{{e}}^{2}H_{k}H^{*}_{k}$, where $H^{*}_{k}$ denotes the adjoint of $H_{k}$. Applying the Hanson-Wright inequality to $\mathcal{A}(e)$, we obtain
$\mathbb{P}\left\{\big{|}\|V_{k}\xi\|^{2}-\mathrm{tr}(\Phi_{v,N}(k))\big{|}>\beta\>\mathrm{tr}(\Phi_{v,N}(k))\right\}\leq 2\exp\left(-c\min\{\beta^{2},\beta\}\frac{\sigma_{{e}}^{4}}{K^{4}}\frac{\mathrm{tr}(\Phi_{v,N}(k))}{\|\Phi_{v,N}(k)\|_{\mathrm{op}}}\right),$ (23)
where we select $\alpha=\beta\mathrm{tr}(\Phi_{v,N}(k))$. We simplify the expression by choosing $c=1/144$ and exploiting the fact that the sub-Gaussian parameter upper-bounds the variance that $K\geq\sigma_{{e}}$.
Step b: covering argument. To control the supremum in (22), we need to discretize the unit sphere. In particular, we consider an $\epsilon-$net and apply the following result.
@@
$\mathbb{P}\left\{\|V^{*}_{k}V_{k}-\mathrm{tr}(\Phi_{v,N}(k))I_{d_{{u}}}\|_{\mathrm{op}}>2\beta\>\mathrm{tr}(\Phi_{v,N}(k))\right\}\leq\\ 9^{2d_{{u}}}2\exp\left(-c\min\left\{\beta^{2},\beta\right\}\frac{\sigma_{{e}}^{4}}{K^{4}}\frac{\mathrm{tr}(\Phi_{v,N}(k))}{\|\Phi_{v,N}(k)\|_{\mathrm{op}}}\right)$ (24)
We only need to simplify the final expression. We invoke the following lemma.
@@
Let $s=\min\{\beta,\beta^{2}\}$, which implies that $\beta=\max\{s,s^{2}\}$. Then, the event
@@
Since $\mathcal{E}_{1}\subset\mathcal{E}_{2}$ the result follows from (24). ∎

### v2

@@ 1 Introduction
Our paper is related to non-parametric system identification, which includes works on both time [20] and frequency domain [21]. Using Gaussian Processes as in [21], where the unknown frequency response follows a Gaussian prior, we can also obtain finite sample guarantees. Here, we follow a different approach and we do not consider Gaussian priors. Note that in this work we focus on qualitative data-independent bounds linking sample requirements to system theoretic properties. Data-dependent bounds, which are arguably more suitable for applications, have also been studied before [22, 23, 24].
Notation. Let $(H,\mathbb{F},\langle\cdot,\cdot\rangle_{H})$ be a Hilbert space with field $\mathbb{F}=$ or $\mathbb{C}$ and inner product $\langle\cdot,\cdot\rangle_{H}$. For any vector $x\in H$, let $\|x\|\triangleq\sqrt{\langle x,x\rangle_{H}}$ denote the inner product norm. Let $H,V$ be Hilbert spaces with $\mathbb{F}=$ or $\mathbb{C}$ and let $\mathcal{A}:H\rightarrow V$ be any linear map. Let $\|\mathcal{A}\|_{\mathrm{op}}\triangleq\sup_{\|x\|=1}\|\mathcal{A}(x)\|$ denote the operator norm and $\mathcal{A}^{*}$ denote the adjoint operator. If $\left\{b_{i}\right\}_{i\in\mathcal{I}}$ is an orthonormal basis, the Hilbert-Schmidt or Frobenius norm is defined as $\|\mathcal{A}\|_{\mathrm{F}}^{2}\triangleq\sum_{i\in\mathcal{I}}\|\mathcal{A}(b_{i})\|^{2}$. Let $\mathcal{L}_{2}(p)\triangleq\{\left\{x_{k}\right\}^{\infty}_{k=1}:x_{k}\in^{p},\,\sum_{k=1}^{\infty}\|x_{k}\|^{2}<\infty\}$ be the Hilbert space of $p$-dimensional square summable sequences. A universal constant is a constant that is independent of the problem at hand, e.g., the system or the algorithm. For any integer $M$, let $[M]\triangleq{0,\dots,M-1}$. The $\mathcal{H}_{\infty}$ norm of $G(e^{j\omega}):[0,2\pi)\rightarrow\mathbb{C}^{d_{1}\times d_{2}}$ is given by $\sup_{\omega\in[0,2\pi)}\|G(e^{j\omega})\|_{\mathrm{op}}$; it is denoted by $\|G\|_{\mathcal{H}_{\infty}}$.
@@ 2 Problem formulation
where $h_{t}\in\mathbb{R}^{d_{{y}}\times d_{{e}}}$ are the unknown filter coefficients. Let $e_{t}\in\mathbb{R}^{d_{{e}}}$ be i.i.d. zero mean, with covariance $\mathbb{E}{e_{t}e_{t}^{\top}}=\sigma_{e}^{2}I_{d_{{e}}}$, and $K^{2}$-sub-Gaussian [4], i.e., for any $\xi\in^{d_{{e}}}$
@@
$\left\|u_{t}\right\|\leq D_{{u}},\,\text{ for all }t\in\mathbb{Z}$
@@ 2.1 Empirical Transfer Function Estimate
The goal of non-parametric frequency domain identification is to estimate the frequency response $G(e^{j\omega})$, given input-output data. We assume access to $d_{{u}}$ experiments of length $N$, that is, data $(u^{(i)}_{0},y^{(i)}_{0},\dots,u^{(i)}_{N-1},y^{(i)}_{N-1})$, for $i=1,\dots,d_{{u}}$. This brings the total number of samples to $N_{\mathrm{tot}}\triangleq d_{{u}}N$. We assume that the trajectories are statistically independent and we leave the single trajectory case for future work.
@@ 2.2 Excitation Method
for some $0<\sigma^{2}_{u,\ell}$.
By definition, for Assumptions 23 to be consistent, we need $\sum_{\ell=0}^{M-1}\sigma^{2}_{u,\ell}\leq MD_{{u}}^{2}$, where $D_{{u}}$ is the input upper bound of Assumption 2.
@@ 2.3 Objective
Fix a frequency resolution $M<N$ such that $M$ divides $N$ and denote their ratio by $N_{p}=N/M$. Consider $d_{{u}}$ independent input-ouput trajectories of length $N$$\{u^{(i)}_{t},y^{(i)}_{t}\}_{t\in[N]}$, for $i=1,\dots,d_{{u}}$, generated by system (1) with excitation inputs as in Assumption 3. Fix a failure probability $10<\delta<$ 1. Determine $\epsilon_{\ell}>0$, $\ell\in[M]$ such that
@@ 3 Finite-sample guarantees for the ETFE
In this section, we focus on estimating the frequency response at $\{2\pi\ell/M\}_{\ell\in[M]}$, that is, the selected frequencies. Following the convention of (4), we define the stacked DFTs of the noises and the noiseless outputs as
@@
The exact value of the universal constant can be found in the proof. The first term of the right-hand side captures the transient error $T_{k,N}U^{-1}_{k}$, while the second one captures the error $V_{k}U^{-1}_{k}$ due to stochastic noise. Recall that the total number of samples is equal to $N_{\mathrm{tot}}=d_{{u}}N$. As we increase the number of samples $N_{\mathrm{tot}}$ while keeping $M$ constant, the former term decays at a faster rate of $1/N_{\mathrm{tot}}$ compared to the latter’s $1/\sqrt{N_{\mathrm{tot}}}$. Hence, the non-asymptotic rate is
@@
The rate is similar to the ones for non-asymptotic parametric identification in time-domain [17]; the optimal rate in that line of work is typically of the order $\sqrt{d}$ for some $d$ scaling with the number of unknown parameters. Here, we have a similar scaling of $\sqrt{M}(\sqrt{d_{{u}}}+\sqrt{d_{{y}}})$ (ignoring log terms) times an additional $\sqrt{d_{{u}}}$ dimensional dependence. This is an artifact of imposing a strict input norm bound in Assumption 2. If we allow $\sigma_{{u}},D_{{u}}$ to scale with $\sqrt{d_{{u}}}$ (as is the case for white-noise inputs in the time-domain [17]), we can remove this extra term.
@@
Lemma 1 (Stochastic Transient [1, Sec. 6.3]).
Denote the power spectrum of the noise at frequency $\omega_{k}=2\pi k/N$, for some $k\in[N]$, by $\Phi_{v}(k)\triangleq\sum_{t=-\infty}^{\infty}R_{t}e^{-j\omega_{k}t}$, with $R_{-t}=R_{t}^{\top}$. We have
$\sup_{k\in[N]}\|\Phi_{v}(k)-\Phi_{v,N}(k)\|_{\mathrm{op}}\leq 2\|R\|_{\star}/N.$
@@ 3.1 Deterministic transient
Lemma 2 (Deterministic Transient [1, Sec. 2.2]).
@@
$\sup_{k\in[N]}\|T_{k,N}\|_{\mathrm{op}}\leq 2\|G\|_{\star}D_{{u}}/\sqrt{N}.$
@@ 3.3 Noise concentration
Fix a $k\in[N]$. Under Assumption 1, for any $s>0$
$\mathbb{P}\left\{\left|\|V_{k}\|_{\mathrm{op}}-\sqrt{\mathrm{tr}(\Phi_{v,N}(k))}\right|>2s\sqrt{\mathrm{tr}(\Phi_{v,N}(k))}\right\}\leq 9^{2d_{{u}}}2\exp\left(-\frac{1}{144}s^{2}\frac{\sigma_{{e}}^{4}}{K^{4}}\frac{\mathrm{tr}(\Phi_{v,N}(k))}{\|\Phi_{v,N}(k)\|_{\mathrm{op}}}\right).$ (14)
@@ 6 Conclusion and Future Work
We provide finite-sample guarantees for the ETFE over a selected frequency grid, in the case of open-loop periodic excitation and under strict stability assumptions. By tuning the frequency resolution and exploiting Lipschitz continuity, we also obtain estimation guarantees in the $\mathcal{H}_{\infty}$ norm. An interesting direction for future work is studying finite-sample non-parametric least squares [5, 16] in the frequency domain. This approach could lead to interesting connections between function class complexity and experiment design. Moreover, adding more structure, beyond Lipschitz continuity, will lead to faster rates. Other topics that are left for future work include the estimation of the noise statistics and extending the guarantees to different variations of the excitation method and the ETFE [28]. Finally, extending minimax lower bounds to this setting is also open [29].
@@ Appendix A Proof of Theorem 2
$\begin{bmatrix}\mathcal{A}_{11}&\mathcal{A}_{12}&\cdots\\ \mathcal{A}_{21}&\mathcal{A}_{22}&\cdots\\ \vdots\\ \mathcal{A}_{d1}&\mathcal{A}_{d2}&\cdots\end{bmatrix},$ (18)
@@
$\|\mathcal{A}\|_{\mathrm{F}}^{2}\triangleq\sum_{i=1}^{d}\sum_{j=1}^{\infty}\|\mathcal{A}_{ij}\|_{\mathrm{F}}^{2}.$
@@
Hence, by the Kolmogorov maximal inequality [30, Th. 2.5.3], the series $\lim_{k\rightarrow\infty}\sum_{j=1}^{k}{\mathcal{A}_{ij}z_{j}}$ converges almost surely for any $i=1,\dots,d$ and, thus, $\mathcal{A}(z)$ is almost surely well-defined. ∎
@@
We will consider finite-dimensional truncations of the quadratic form $\|\mathcal{A}(z)\|^{2}$, apply the result of [17], and take limits. Let $\hat{\mathcal{A}}^{(k)}:\mathcal{L}_{2}(p)\rightarrow^{d}$ be the truncated map with semi-infinite matrix representation
@@
Step a): truncation. First, we will establish the desired inequality for all $X_{k}$. Notice that $\|\hat{\mathcal{A}}^{(k)}(z)\|^{2}$ is a finite dimensional quadratic form. Hence, Proposition A.1 of [17] applies and provides the bound
@@
$|\|\mathcal{A}(z)\|^{2}-\|\hat{\mathcal{A}}^{(k)}(z)\|^{2}|\leq 2\left|\left(\hat{\mathcal{A}}^{(k)}(z)\right)^{\top}(\mathcal{A}-\hat{\mathcal{A}}^{k})(z)\right|+\|(\mathcal{A}-\hat{\mathcal{A}}^{k})(z)\|^{2}.$
Let $C_{z}=\sup_{t\geq 1}\mathbb{E}z^{\top}_{t}z_{t}$; due to sub-Gaussianity $C_{z}=cpK^{2}$ for some universal constant $c$–see proof of Lemma 5. Invoking the Cauchy-Schwarz inequality
$\mathbb{E}|\|\mathcal{A}(z)\|^{2}-\|\hat{\mathcal{A}}^{(k)}(z)\|^{2}|\leq 2C_{z}\|\mathcal{A}\|_{\mathrm{F}}\|\mathcal{A}-\hat{A}^{(k)}\|_{\mathrm{F}}+C_{z}\|\mathcal{A}-\hat{A}^{(k)}\|_{\mathrm{F}}^{2}\rightarrow 0,$
@@
The proof of Theorem 2 now follows from Lemma 6 and the Chernoff bound method. The remaining details are omitted since they are identical to [17, proof of Th. 2.1 and Th. B.1].
@@ Appendix B Proof of Lemma 4
First, we will establish a high probability bound on $V_{k}V^{*}_{K}$. Then we we use that to obtain a high probability bound on the operator norm $\|V_{k}\|_{\mathrm{op}}$.
Step 1): Bound on $V_{k}V^{*}_{k}$. Due to the independence between the experiments
$\mathbb{E}V^{*}_{k}V_{k}=\mathrm{diag}(\mathbb{E}(V^{(i)})^{*}_{k}V^{(i)}_{k})=\mathrm{tr}(\Phi_{v,N}(k))I_{d_{{u}}}.$ (23)
@@
Step 1a): fixed $\xi$. Observe that $V^{(i)}_{k}=\mathcal{F}_{k}((h*e^{(i)}))$ is the output of a linear map applied to the infinite sequence $e^{(i)}\triangleq\{e^{(i)}_{N-1},e^{(i)}_{N-2},\dots\}$, where * denotes convolution. We can rewrite the above linear relation using the notation $V^{(i)}_{k}=H_{k}(e^{(i)})$, where $H_{k}$ represents the map $\mathcal{F}_{k}\circ((h*\cdot))$. For convenience, let us group all noise sequences $e^{(i)}$ into one single sequence $e$. We use the enumeration
@@
$\Xi=\begin{bmatrix}\xi^{*}\otimes I_{d_{{e}}}&0&0&\cdots\\ 0&\xi^{*}\otimes I_{d_{{e}}}&0&\cdots\\ 0&0&\xi^{*}\otimes I_{d_{{e}}}\\ \vdots&&&\ddots\end{bmatrix},$
where $\otimes$ denotes the Kronecker product. Since $\xi^{*}\xi=1$, $\Xi\circ\Xi^{*}$ is the identity operator and has unit norm. Note that $(H_{k}\circ\Xi)(e)\in\mathbb{C}^{d_{{y}}}$ is complex-valued, while the Hanson-Wright inequality applies to real valued maps. To resolve this issue, we embed $V_{k}\xi$ to a real vector space. Let $\Re(H_{k}\circ\Xi)$ denote the real part of the operator $H_{k}\circ\Xi$ with $\Im(H_{k}\circ\Xi)$ the respective imaginary part. Define the lifted operator
$\mathcal{A}(e)\triangleq\begin{bmatrix}\Re(H_{k}\circ\Xi)(e)\\ \Im(H_{k}\circ\Xi)(e)\end{bmatrix}\in^{2d_{{y}}}.$
Our goal is apply the Hanson-Wright inequality to $\mathcal{A}(e)$. Observe that we have
$\|\mathcal{A}(e)\|^{2}=\|\Re(H_{k}\circ\Xi)(e)\|^{2}+\|\Im(H_{k}\circ\Xi)(e)\|^{2}=\xi^{*}V^{*}_{k}V_{k}\xi=\|V_{k}\xi\|^{2}.$ (25)
Using the representation of $\mathcal{A}$ as a semi-infinite matrix as in (18) and since the noise sequence $e=\{e_{t}\}_{t=1}^{\infty}$ is i.i.d. with covariance $\sigma_{{e}}^{2}I_{d_{{e}}}$, we have
$\mathbb{E}\|\mathcal{A}(e)\|^{2}=\mathbb{E}\sum_{i=1}^{2d_{{y}}}(\sum_{j=1}^{\infty}\mathcal{A}_{ij}e_{j})^{2}=\sigma_{{e}}^{2}\sum_{i=1}^{2d_{{y}}}\sum_{j=1}^{\infty}\|\mathcal{A}_{ij}\|_{\mathrm{F}}^{2}=\sigma_{{e}}^{2}\|\mathcal{A}\|_{\mathrm{F}}^{2}.$
Meanwhile, we have
[…]
Combining the above equalities results in
$\mathbb{E}\|\mathcal{A}(e)\|^{2}=\mathrm{tr}(\Phi_{v,N}(k)),\,\|\mathcal{A}\|_{\mathrm{F}}^{2}=\mathrm{tr}(\Phi_{v,N}(k))\sigma_{{e}}^{-2}.$ (26)
Next, we upper-bound the operator norm of $\mathcal{A}$. Since the sequence $e$ is real valued, we have
$\|\mathcal{A}\|_{\mathrm{op}}=\|H_{k}\circ\Xi\|_{\mathrm{op}}\leq\|H_{k}\|_{\mathrm{op}},$
where the inequality follows from the fact that $\Xi$ has unit norm. Finally, since $V^{(i)}_{k}=H_{k}e^{(i)}$, for any $i=1,\dots,d_{{u}}$, we have
$\Phi_{v,N}(k)=\mathbb{E}V^{(i)}_{k}(V^{(i)})^{*}_{k}=\sigma_{{e}}^{2}H_{k}H^{*}_{k}.$
Combing the above we obtain
$\|\mathcal{A}\|_{\mathrm{op}}\leq\sqrt{\|\Phi_{v,N}(k)\|_{\mathrm{op}}}/\sigma_{{e}}.$ (27)
With (25), (26), (27) at hand we can finally apply the Hanson-Wright inequality to $\mathcal{A}(e)$. Invoking Theorem 2 with $\alpha=\beta\sigma_{{e}}$, we obtain
$\mathbb{P}\left\{\big{|}\|V_{k}\xi\|^{2}-\mathrm{tr}(\Phi_{v,N}(k))\big{|}>\beta\>\mathrm{tr}(\Phi_{v,N}(k))\right\}\leq 2\exp\left(-144^{-1}\min\{\beta^{2},\beta\}\frac{\sigma_{{e}}^{4}}{K^{4}}\frac{\mathrm{tr}(\Phi_{v,N}(k))}{\|\Phi_{v,N}(k)\|_{\mathrm{op}}}\right),$ (28)
Note that we simplified the expression by using $1/144\leq 1/(16\sqrt{2})$ and by exploiting the fact that the sub-Gaussian parameter upper-bounds the variance, that is, $K\geq\sigma_{{e}}$.
Step 1b): covering argument. To control the supremum in (24), we need to discretize the unit sphere. In particular, we consider an $\epsilon-$net and apply the following result.
@@
$\mathbb{P}\left\{\|V^{*}_{k}V_{k}-\mathrm{tr}(\Phi_{v,N}(k))I_{d_{{u}}}\|_{\mathrm{op}}>2\beta\>\mathrm{tr}(\Phi_{v,N}(k))\right\}\leq\\ 9^{2d_{{u}}}2\exp\left(-144^{-1}\min\left\{\beta^{2},\beta\right\}\frac{\sigma_{{e}}^{4}}{K^{4}}\frac{\mathrm{tr}(\Phi_{v,N}(k))}{\|\Phi_{v,N}(k)\|_{\mathrm{op}}}\right).$ (29)
Step 2): Bound on the operator norm $\|V_{k}\|_{\mathrm{op}}$. We invoke the following lemma.
@@
Let $s^{2}=\min\{\beta,\beta^{2}\},s\geq 0$, which is equivalent to $\beta=\max\{s,s^{2}\}$. Then, the event
@@
Since $\mathcal{E}_{1}\subset\mathcal{E}_{2}$ the result follows from (29), where we replace $\min\{\beta^{2},\beta\}$ with $s^{2}$. ∎